OpenStudy (anonymous):

Using complete sentences, explain how to graph y = negative one fifthx - 2 using the slope-intercept method.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$y=\frac{1}{5}x-2$$

OpenStudy (whpalmer4):

Remember, in slope-intercept form, y = mx + b, m is the slope, and b is the y-intercept, the point at which the line crosses the y-axis. Slope is the ratio of change in y to change in x. A slope of -1/5 means that for y to change by -1, x has to change by 5. This is a line that is nearly horizontal, sloping down gradually to the right.

OpenStudy (anonymous):

Thank you, but how do I know where to graph it?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Well, you know 1 point: (0, y-intercept) = (0, -2). So the line goes through that point, and it slopes down and to the right, dropping 1 unit on the y-axis for every 5 units moved to the right on the x-axis. Another point would be (0+5, -2-1) = (5, -3). Draw your line through those points. Or, rearrange your equation into standard form, which has x and y on the left side, and a number on the right side: \[Ax + By = C\]which in this case is\[y - \frac{1}{5}x = -2\] (I subtracted x/5 from both sides) or more conveniently after multiplying both sides by 5\[5y - 1x = -10\]The nice thing about standard form is it makes drawing the line easy. The x-intercept (y = 0) is (C/A, 0) and the y-intercept is (0, C/B), so you just plot those points and draw a straight line through them.

OpenStudy (whpalmer4):

For your formula, A = -1, B = 5, C = -10: \[-1x + 5y = -10\] (I should have written in that order, sorry)

OpenStudy (anonymous):

I thought it was something like, m = y1 - y2 over x1 - x2

OpenStudy (whpalmer4):

Yes, that is a formula for determining the slope between two points.

OpenStudy (anonymous):

\[m =\frac{ y1 - y2 }{ x1 - x2 }\]

OpenStudy (whpalmer4):

But we already know the slope, because our equation was in slope-intercept form: y = mx + b

OpenStudy (anonymous):

Ok then for my answer to: Using complete sentences, explain how to graph y = -1/5x - 2 using the slope-intercept method. What should I type in the box?

OpenStudy (whpalmer4):

Well, as I said, you identify the slope from the equation, and the y-intercept to fix a point on the line. Then you draw a line through that point with the slope you found.

OpenStudy (whpalmer4):

Remember, the y-intercept is the point at which the line crosses the y-axis, meaning x = 0, so you'll know that (0, b) is a point on the line. y = mx + b, y = m(0) + b, y = b at the y-intercept.

OpenStudy (anonymous):

Im so confused :S Sorry I just dont get it

OpenStudy (anonymous):

Will you just tell me what you would put in the answer box if you were taking the quiz.

OpenStudy (whpalmer4):

Trust me, you do need to understand this unless this is the very last math quiz you'll ever take :-) Which part don't you understand?

OpenStudy (anonymous):

I understand everything except how to put it on the graph.

OpenStudy (whpalmer4):

What do you need to know to put it on the graph?

OpenStudy (anonymous):

x adn y intercepts, which is -1/5 and -2

OpenStudy (whpalmer4):

Well, that's one way to do it. You know the y-intercept already. You could find the x-intercept using the equation for the line and the knowledge that the x-intercept is where y = 0. But my interpretation of the question is that they want you to use the y-intercept and the slope to graph it. You know the y-intercept. You know the slope. You just need to know what the slope means when you draw the line. I'm guessing you don't have a feel for what the slope means in terms of a graph. Here are some examples: slope of 1 (m = 1) gives a graph that goes up at a 45 degree angle. Every 1 step to the right, you go up 1 step. slope of =1 (m = -1) gives a similar graph, but it goes down at a 45 degree angle. Every 1 step to the right, you go down 1 step. If you draw two lines, one with slope 1, and another with slope -1, you get a nice symmetrical X on your graph. The definition of perpendicular lines is that the product of their slopes = -1. A slope which is > 1 means that the line goes up faster than it goes to the right. A slope of 2, for example, goes up by 2 steps for every step to the right. A slope which is < 1 (but > 0) means that the line goes up slower than it goes to the right. A slope of 1/2 goes up 1 step for every 2 steps to the right. A negative slope behaves like the positive slope would, except that it goes in the opposite direction. A slope of -2 goes down faster than it does to the right, going down 2 steps for every step to the right. A slope of -1/2 goes down 1 step for every 2 steps to the right.

OpenStudy (whpalmer4):

So, plot the point for the y-intercept, then move over and up/down the appropriate number of steps for the slope, plot that point, and draw a straight line through them.

OpenStudy (anonymous):

Ok thanks for you help but I think Im more confused now than I was before. Guess Ill just try and figure it out myself :S